OpenStudy (anonymous):

Using imaginary numbers: solve for the absolute value of the following complex numbers: 6-8i

OpenStudy (anonymous):

This right? \[\left| 6-8i \right|\]

OpenStudy (anonymous):

i guess, i did not understand it during algebra. 6-8i is exactly what it says on my worksheet.

OpenStudy (anonymous):

Because there's no need to use imaginary numbers. It's just this: \[\sqrt{(6)^2+(-8)^2}\] \[=\sqrt{100}\]

OpenStudy (anonymous):

It's using Pythagoras thereom.

OpenStudy (anonymous):

I can draw it for you to better understand it.

OpenStudy (anonymous):

oh, ok. so just square everything and put it all under a square root?

OpenStudy (anonymous):

and then whatever the square root is is your answer?

OpenStudy (anonymous):